Circular No. 5905

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CASSIOPEIAE 1993 B. Schmidt, Harvard-Smithsonian Center for Astrophysics, reports

that a spectrogram (range 380-900 nm) taken on Dec. 13.3 UT with the Multiple Mirror Telescope confirms that this object (cf. IAUC 5902, 5904) is indeed an iron-rich nova. Hα and Hβ are very strong in emission; Hγ is also seen. A multitude of lines exist, the strongest ones including those at 589.3 nm (Na I D); 777.3 and 844.6 nm (O I); 850.0 and 854.6 nm (Ca II); and many Fe II lines (442.0, 458.9, 463.2, 467.1, 473.4, 492.4, 501.8, 516.9, 519.7, 523.7, 527.8, 531.8, 536.6, 616.1, and 625.0 nm).

Photoelectric photometry: Dec. 12.70 UT, V = 6.39 ± 0.02, B–V = +0.98 ± 0.0 (H. Mikuž, Ljubljana, Slovenia; comparison star HR 9010); 13.1, V = 6.461, b–y (Strömgren) = +0.557 (B. Skiff, Lowell Observatory; based on V = 6.246 and b–y = +0.630 for HR 8985, and on V = 7.213, b–y = +0.112 for HD 222514).

Visual magnitude estimates: Dec. 13.70 UT, 6.4 (L. Kiss, Szeged, Hungary); 14.25, 6.4 (G. M. Hurst, Basingstoke, U.K.).

PSR 1706–44 T. Kifune, N. Hayashida, T. Tamura, and M. Teshima, Institute of

Cosmic Ray Research, University of Tokyo; T. Tanimori, S. Ogio, F. Kakimoto, T. Tsukagoshi, and T. Yoshikoshi, Department of Physics, Tokyo Institute of Technology; Y. Matsubara and Y. Muraki, Solar and Terrestrial Environmental Laboratory, Nagoya University; Y. Mizumoto and T. Suda, Department of Physics, Kobe University; T. Hara, Yuge National College of Maritime Technology; H. Fujii, S. Kabe, and Y. Watase, National Laboratory for High Energy Physics (KEK), Japan; M. Fujimoto, National Astronomical Observatory, Japan; and P. G. Edwards, J. R. Patterson, G. P. Rowell, M. D. Roberts, and G. J. Thornton, Department of Physics, University of Adelaide, communicate: “Photons of energy 1 TeV have been detected from the Gamma Ray Observatory source PSR 1706 − 44. Observations were made using the 3.8-m imaging telescope of Project CANGAROO which is located at Woomera, South Australia. After cuts, a clear peak was seen in the alpha distribution (see Punch et al. 1992, Nature 358, 477). The signal was seen in July and August 1992 and 1993 with a combined significance as large as 10σ. The excess corresponds to a constant flux of ∼ 10^{−11} cm^{−2} s^{−1} above the 1-TeV telescope threshold.”

1993 December 14 Daniel W. E. Green